# Are Graph Neural Networks Optimal Approximation Algorithms?

### v5

@@ 5 Conclusion
We have presented OptGNN, a GNN that can capture provably optimal message passing algorithms for a large class of combinatorial optimization problems. OptGNN achieves the appealing combination of obtaining approximation guarantees while also being able to adapt to the data to achieve improved results. Empirically, we observed that the OptGNN architecture achieves strong performance on a wide range of datasets and on multiple problems. OptGNN opens up several directions for future exploration, such as the design of more powerful and sound (neural) rounding procedures that can secure approximation guarantees, the construction of neural certificates that improve upon the ones we described in Appendix C.2, and the design of neural SDP-based branch and bound solvers.
6 Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@ Appendix B Max-3-SAT OptGNN
$\text{MESSAGE}[v_{ij}\rightarrow v_{k}]=\sum_{c\in\mathcal{P}:(x_{i},x_{j},x_{k})\in c}\Biggr{[}-\frac{1}{8}\Bigg{(}-\tau_{i}\tau_{j}\tau_{k}\frac{1}{3}v_{k}\Bigg{)}+2\rho\left[(\langle v_{k},v_{ij}\rangle-\langle v_{i},v_{jk}\rangle)v_{k}+(\langle v_{k},v_{ij}\rangle-\langle v_{j},v_{ik}\rangle)v_{k}\right]\Biggr{]}$ (47)
@@
$\text{MESSAGE}[v_{i}\rightarrow v_{j}]=\sum_{c\in\mathcal{P}:(x_{i},x_{j},x_{k})\in c}\Biggr{[}-\frac{1}{8}\Bigg{(}-\tau_{i}\tau_{j}\frac{1}{2}v_{i}\Bigg{)}+2\rho\left[(\langle v_{i},v_{j}\rangle-\langle v_{ij},v_{\emptyset}\rangle)v_{i}+(\langle v_{ik},v_{jk}\rangle-\langle v_{i},v_{j}\rangle)(-v_{i})\right]\Biggr{]}$ (48)
@@ Appendix C Optimality of Message Passing for Max-CSP
Our primary theoretical result is that a polynomial time message passing algorithm on an appropriately defined constraint graph computes the approximate optimum of SDP 1 which is notable for being an SDP that achieves the Unique Games optimal integrality gap.
Our proof roadmap is simple. First, we design an SDP relaxation SDP 1 for Max-k-CSP that is provably equivalent to the SDP of Raghavendra (2008) and therefore inherits its complexity theoretic optimality. Finally, we design a message passing algorithm to approximately solve SDP 1 in polynomial time to polynomial precision. Our message passing algorithm has the advantage of being formulated on an appropriately defined constraint graph. For a Max-k-CSP instance $\Lambda$ with $N$ variables, $|\mathcal{P}|$ predicates, over an alphabet of size $q$, it takes $|\mathcal{P}|q^{k}$ space to represent the Max-CSP. To declutter notation, we let $\Phi$ be the size of the Max-CSP which is equal to $|\mathcal{P}|q^{k}$ Our message passing algorithm achieves an additive $\epsilon$ approximation in time $poly(\epsilon^{-1},\Phi,\log(\delta^{-1}))$ which is then polynomial in the size of the CSP and inverse polynomial in the precision.
@@
There is an SDP relaxation of equation 55 that is the ”qualitatively most powerful assuming the Unique Games conjecture” Raghavendra (2008). More specifically, the integrality gap of the SDP achieves the Unique Games optimal approximation ratio. Furthermore, there exists a rounding that achieves its integrality gap.
@@
Specifically, let $z$ be a subset of the CSP variables $z\subset\{x_{i}\}_{i\in[N]}$ of size $k$. Let $X_{z}$ denote the multilinear moment $X_{z}\vcentcolon=\prod_{i\in z}x_{i}$. Then $\tilde{\mathbb{E}}_{\mu}[X_{z}]$ denotes the SDP variable corresponding to the multilinear moment $\mathbb{E}_{\mu}[X_{z}]$. Of course optimizing over the space of distributions $\mu$ over $\mathcal{V}$ is intractable, and so we opt for optimizing over the space of low degree pseudodistributions and their associated pseudoexpecation functionals. See Barak & Steurer (2014) for references therein.
@@
$\min_{x_{1},x_{2},\dots,x_{N}}$ $\frac{1}{|\mathcal{P}|}\sum_{P_{z}\subset\mathcal{P}}\tilde{\mathbb{E}}_{\mu}[-P_{z}(X_{z})]\vcentcolon=\sum_{P_{z}\in\mathcal{P}}\sum_{s\subseteq z}w_{s}\frac{1}{|\mathcal{C}(s)|}\sum_{g,g^{\prime}\subseteq s:\zeta(g,g^{\prime})=s}\langle v_{g},v_{g^{\prime}}\rangle$ (60)
@@
For Max-k-CSP instance $\Lambda$, The SDP of SDP 1 (top of pg.19) is at least as tight as the SDP of Raghavendra (2008).
@@
The SDP of Raghavendra (2008) is a degree $2$ SoS SDP augmented with $k$-local distributions for every predicate $P\in\mathcal{P}$. By using the vectors of the cholesky decomposition and constraining them to be unit vectors we automatically capture degree $2$ SoS. To capture $k$ local distributions we simply enforce degree $2k$ SoS on the boolean hypercube for the domain of every predicate. This can be done with the standard vector formulation written in SDP 1 (top of pg.18). See Barak & Steurer (2014) for background and references. ∎
@@
Noticing that the form of the gradient depends only on the vectors in the neighborhood of the constraint graph $G_{\Lambda}$ we arrive at our message passing algorithm. The key to our proof is bounding the number of iterations required to optimize equation 63 to sufficient accuracy to be an approximate global optimum of SDP 1.
@@
3:  $\eta,\psi,\sigma\leftarrow n^{-100}$             {Initialize step size, noise threshold, and noise variance}
@@
$+2\rho\Bigg{[}\sum_{\begin{subarray}{c}P_{z}\in\mathcal{P}\\ \text{s.t }w\subseteq z\end{subarray}}\sum_{\begin{subarray}{c}w^{\prime},h,h^{\prime}\subseteq s\\ \text{s.t }\zeta(w,w^{\prime})=\zeta(h,h^{\prime})\end{subarray}}\big{(}\langle v_{w}^{t},v_{w^{\prime}}^{t}\rangle-\langle v_{h}^{t},v_{h^{\prime}}^{t}\rangle\big{)}v_{w}^{\prime t}+(\|v_{w}^{t}\|^{2}-1)v_{w}^{t}\Bigg{]}\Biggr{]}$ (76)
@@
17:  Output: vectors corresponding to solution to SDP 1 (top of pg.18)
@@
Algorithm 2 computes in $poly(\epsilon^{-1},\Phi,\log(\delta^{-1}))$ iterations a set of vectors $\mathbf{v}\vcentcolon=\{\hat{v}_{s}\}$ for all $s\subseteq\mathcal{S}(P)$ for all $P\in\mathcal{P}$ that satisfy the constraints of SDP 1 to error $\epsilon$ and approximates the optimum of SDP 1 to error $\epsilon$ with probability $1-\delta$
[…]
@@
Thus far we have established the form of the gradient. To prove the gradient iteration converges we reference the literature on convergence of perturbed gradient descent (Jin et al., 2017) which we rewrite in Theorem F.1. First we note that the SDP 1 has $\ell$ smooth gradient for $\ell\leq poly(\rho,B)$ and has $\gamma$ lipschitz Hessian for $\gamma=poly(\rho,B)$ which we arrive at in Lemma E.2 and Lemma F.1. The proofs of Lemma E.2 and Lemma F.1 are technically involved and form the core hurdle in arriving at our proof.
Then by Theorem F.1 the iteration converges to an $(\epsilon^{\prime},\gamma^{2})$-SOSP Definition Definition in no more than $\tilde{O}(\frac{1}{\epsilon^{\prime 2}})$ iterations with probability $1-\delta$. Now that we’ve established that the basic gradient iteration converges to approximate SOSP, we need to then prove that the approximate SOSP are approximate global optimum. We achieve this by using the result of Bhojanapalli et al. (2018) lemma 3 which we restate in Lemma C.2 for convenience.
@@
Let the global optimum of SDP 1 be denoted $\tilde{\mathbb{E}}_{\tilde{\mu}}[\tilde{\mathbf{x}}]$ with a cholesky decomposition $\tilde{\mathbf{v}}$. Let $\hat{\mathbf{v}}$ be the set of vectors outputted by Algorithm 2 with associated pseudoexpectation $\tilde{\mathbb{E}}_{\hat{\mu}}[\hat{\mathbf{x}}]$. Then, we can bound
@@
Here the first equality is by definition, and the inequality is by the convexity of $\mathcal{H}_{\rho}$. Moving on, we use the fact that the min eigenvalue of the hessian of overparameterized loss $\nabla^{2}\mathcal{L}_{\rho}(\hat{\mathbf{v}})\succeq-\gamma\sqrt{\epsilon^{\prime}}$ implies the min eigenvalue of the gradient of the convex loss $\lambda_{\text{min}}(\nabla\mathcal{H}_{\rho}(\tilde{\mathbb{E}}[\hat{\mathbf{x}}]))\geq-\gamma\sqrt{\epsilon^{\prime}}$. This fact is invoked in Bhojanapalli et al. (2018) lemma 3 which we adapt to our setting in Lemma C.2. Subsequently, we adapt the lines of their argument in Lemma C.2 most relevant to our analysis which we detail here for the sake of completeness.
@@
Here the first inequality follows by a standard inequality of frobenius inner product, the second inequality follows by Cauchy-Schwarz, the third inequality follows by the $(\epsilon^{\prime},\gamma^{2})$-SOSP conditions on both the min eigenvalue of the hessian and the norm of the gradient, the final two inequalities follow from knowing the main diagonal of $\tilde{\mathbb{E}}[\hat{\mathbf{x}}]$ is the identity and that every vector in $\tilde{\mathbf{v}}$ is a unit vector up to inverse polynomial error $poly(\rho^{-1},2^{k})$. For this last point see the proof in Lemma E.1. Therefore if we set $\epsilon^{\prime}=poly(\epsilon,2^{-k})$ we arrive at any $\epsilon$ error. Therefore we have established our estimate $\hat{v}$ is approximates the global optimum of the quadratically penalized objective i.e $\mathcal{H}_{\rho}(\tilde{\mathbb{E}}[\hat{\mathbf{x}}])-\mathcal{H}_{\rho}(\tilde{\mathbb{E}}[\tilde{\mathbf{x}}])\leq\epsilon$. To finish our proof, we have to bound the distance between the global optimum of the quadratically penalized objective $\mathcal{H}_{\rho}(\tilde{\mathbb{E}}[\tilde{\mathbf{x}}])$ and SDP$(\Lambda)$ the optimum of SDP 1. This is established for $\rho$ a sufficiently large $poly(\epsilon^{-1},2^{k})$ in Lemma E.1. This concludes our proof that the iterates of Algorithm 2 converge to the solution of SDP 1. ∎
In our proof above, the bound on approximate SOSP being approximate global optimum is built on the result of Bhojanapalli et al. (2018) which we rephrase for our setting below.
@@
[Bhojanapalli et al. (2018) lemma 3 rephrased] Let $\mathcal{L}_{\rho}(\cdot)$ be defined as in equation 77 and let $\mathcal{H}_{\rho}(\cdot)$ be defined as in equation 85. Let $U\in\mathbb{R}^{\Phi\times\Phi}$ be an $(\epsilon,\gamma)$-SOSP of $\mathcal{L}_{\rho}(\cdot)$, then
@@
The following Lemma E.1 establishes that for a sufficiently large penalty parameter $\rho=poly(\epsilon^{-1},2^{k})$ the optimum of the penalized problem and the exact solution to SDP 1 are close.
@@
Let $\Lambda$ be a Max-k-CSP instance, and let SDP$(\Lambda)$ be the optimum of SDP 1. Let $\mathcal{L}_{\rho}(\mathbf{v})$ be the quadratically penalized objective
@@
For an objective matrix $C$ and constraint matrices $\{A_{i}\}_{i\in\mathcal{F}}$ in some constraint set $\mathcal{F}$. We will invoke specific properties of SDP 1 to enable our analysis. First we define the penalized objective in this generic form
@@
We upper bound the RHS using the robustness theorem of Raghavendra & Steurer (2009a) restated in the appendix Theorem F.2 which states that an SDP solution that violates the constraints by a small perturbation changes the objective by a small amount. Thus we obtain,
@@
SDP 2 SDP Vector Formulation for Max-k-CSP General Alphabet (Equivalent to UGC optimal)
@@
C.1 Definitions
Definition (Message passing algorithm).
Given a graph $G=(V,E)$ with a set of vectors $\mathbf{v}=\{v_{i}\}_{i\in V}$ a message passing algorithm for $T$ iterations is an update of the form; for vector $v_{i}$ in $\mathbf{v}$ and for iteration $t\in[T]$,
$v_{i}^{t+1}=\text{UPDATE}(\{v_{j}^{t}\}_{j\in N(i)},v_{i}^{t})$
For an arbitrary polynomial time computable function $\text{UPDATE}:\mathbb{R}^{r|N(i)|}\times\mathbb{R}^{r}\rightarrow\mathbb{R}^{r}$
C.2 Neural Certification Scheme

### v7

@@ 5 Conclusion
We presented OptGNN, a GNN that can capture provably optimal message passing algorithms for a large class of combinatorial optimization problems. OptGNN achieves the appealing combination of obtaining approximation guarantees while also being able to adapt to the data to achieve improved results. Empirically, we observed that the OptGNN architecture achieves strong performance on a wide range of datasets and on multiple problems. OptGNN opens up several directions for future exploration, such as the design of powerful rounding procedures that can secure approximation guarantees, the construction of neural certificates that improve upon the ones we described in Appendix C.1, and the design of neural SDP-based branch and bound solvers.
6 Acknowledgment
The authors would like to thank Ankur Moitra, Sirui Li, and Zhongxia Yan for insightful discussions in the preparation of this work. Nikolaos Karalias is funded by the SNSF, in the context of the project ”General neural solvers for combinatorial optimization and algorithmic reasoning” (SNSF grant number: P500PT_217999). Stefanie Jegelka and Nikolaos Karalias acknowledge support from NSF AI Institute TILOS (NSF CCF-2112665). Stefanie Jegelka acknowledges support from NSF award 2134108.
@@ Appendix B Max-3-SAT OptGNN
$\text{MESSAGE}[v_{ij}\rightarrow v_{k}]=\sum_{c\in\mathcal{P}:(x_{i},x_{j},x_{k})\in c}\Biggr{[}-\frac{1}{8}\Bigg{(}-\tau_{i}\tau_{j}\tau_{k}\frac{1}{3}v_{k}\Bigg{)}$ (44)
$+2\rho\left[(\langle v_{k},v_{ij}\rangle-\langle v_{i},v_{jk}\rangle)v_{k}+(\langle v_{k},v_{ij}\rangle-\langle v_{j},v_{ik}\rangle)v_{k}\right]\Biggr{]}$ (45)
@@
$\text{MESSAGE}[v_{i}\rightarrow v_{j}]=\sum_{c\in\mathcal{P}:(x_{i},x_{j},x_{k})\in c}\Biggr{[}-\frac{1}{8}\Bigg{(}-\tau_{i}\tau_{j}\frac{1}{2}v_{i}\Bigg{)}$ (46)
$+2\rho\left[(\langle v_{i},v_{j}\rangle-\langle v_{ij},v_{\emptyset}\rangle)v_{i}+(\langle v_{ik},v_{jk}\rangle-\langle v_{i},v_{j}\rangle)(-v_{i})\right]\Biggr{]}$ (47)
@@ Appendix C Optimality of Message Passing for Max-CSP
Our primary theoretical result is that a polynomial time message passing algorithm on an appropriately defined constraint graph computes the approximate optimum of SDP 2 which is notable for being an SDP that achieves the Unique Games optimal integrality gap.
Our proof roadmap is simple. First, we design an SDP relaxation SDP 2 for Max-k-CSP that is provably equivalent to the SDP of raghavendra2008optimal and therefore inherits its complexity theoretic optimality. Finally, we design a message passing algorithm to approximately solve SDP 2 in polynomial time to polynomial precision. Our message passing algorithm has the advantage of being formulated on an appropriately defined constraint graph. For a Max-k-CSP instance $\Lambda$ with $N$ variables, $|\mathcal{P}|$ predicates, over an alphabet of size $q$, it takes $|\mathcal{P}|q^{k}$ space to represent the Max-CSP. To declutter notation, we let $\Phi$ be the size of the Max-CSP which is equal to $|\mathcal{P}|q^{k}$ Our message passing algorithm achieves an additive $\epsilon$ approximation in time $poly(\epsilon^{-1},\Phi,\log(\delta^{-1}))$ which is then polynomial in the size of the CSP and inverse polynomial in the precision.
@@
There is an SDP relaxation of equation 52 that is the ”qualitatively most powerful assuming the Unique Games conjecture” raghavendra2008optimal. More specifically, the integrality gap of the SDP achieves the Unique Games optimal approximation ratio. Furthermore, there exists a rounding that achieves its integrality gap.
@@
Specifically, let $z$ be a subset of the CSP variables $z\subset\{x_{i}\}_{i\in[N]}$ of size $k$. Let $X_{z}$ denote the multilinear moment $X_{z}\vcentcolon=\prod_{i\in z}x_{i}$. Then $\tilde{\mathbb{E}}_{\mu}[X_{z}]$ denotes the SDP variable corresponding to the multilinear moment $\mathbb{E}_{\mu}[X_{z}]$. Of course optimizing over the space of distributions $\mu$ over $\mathcal{V}$ is intractable, and so we opt for optimizing over the space of low degree pseudodistributions and their associated pseudoexpecation functionals. See barak2014sum for references therein.
@@
$\min$ $\frac{1}{|\mathcal{P}|}\sum_{P_{z}\subset\mathcal{P}}\tilde{\mathbb{E}}_{\mu}[-P_{z}(X_{z})]\vcentcolon=\sum_{P_{z}\in\mathcal{P}}\sum_{s\subseteq z}w_{s}\frac{1}{|\mathcal{C}(s)|}\sum_{g,g^{\prime}\subseteq s:\zeta(g,g^{\prime})=s}\langle v_{g},v_{g^{\prime}}\rangle$ (57)
@@
For Max-k-CSP instance $\Lambda$, The SDP of SDP 2 is at least as tight as the SDP of raghavendra2008optimal.
@@
The SDP of raghavendra2008optimal is a degree $2$ SoS SDP augmented with $k$-local distributions for every predicate $P\in\mathcal{P}$. By using the vectors of the cholesky decomposition and constraining them to be unit vectors we automatically capture degree $2$ SoS. To capture $k$ local distributions we simply enforce degree $2k$ SoS on the boolean hypercube for the domain of every predicate. This can be done with the standard vector formulation written in SDP 2. See barak2014sum for background and references. ∎
@@
Noticing that the form of the gradient depends only on the vectors in the neighborhood of the constraint graph $G_{\Lambda}$ we arrive at our message passing algorithm. The key to our proof is bounding the number of iterations required to optimize equation 60 to sufficient accuracy to be an approximate global optimum of SDP 2.
@@
3:  $\eta,\psi,\sigma\leftarrow n^{-2}$             {Initialize step size, noise threshold, and noise variance}
@@
$+2\rho\Bigg{[}\sum_{\begin{subarray}{c}P_{z}\in\mathcal{P}\\ \text{s.t }w\subseteq z\end{subarray}}\sum_{\begin{subarray}{c}w^{\prime},h,h^{\prime}\subseteq s\\ \text{s.t }\zeta(w,w^{\prime})=\zeta(h,h^{\prime})\end{subarray}}\big{(}\langle v_{w}^{t},v_{w^{\prime}}^{t}\rangle-\langle v_{h}^{t},v_{h^{\prime}}^{t}\rangle\big{)}v_{w}^{\prime t}+(\|v_{w}^{t}\|^{2}-1)v_{w}^{t}\Bigg{]}\Biggr{]}$
@@
17:  Output: vectors corresponding to solution to SDP 2
@@
Algorithm 3 computes in $O(\epsilon^{-4}\Phi^{4}\log(\delta^{-1}))$ iterations a set of vectors $\mathbf{v}\vcentcolon=\{\hat{v}_{s}\}$ for all $s\subseteq\mathcal{S}(P)$ for all $P\in\mathcal{P}$ that satisfy the constraints of SDP 2 to error $\epsilon$ and approximates the optimum of SDP 2 to error $\epsilon$ with probability $1-\delta$
where $\text{SDP}(\Lambda)$ is the optimum of SDP 2.
@@
Thus far we have established the form of the gradient. To prove the gradient iteration converges we reference the literature on convergence of perturbed gradient descent (jin2017escape) which we rewrite in Theorem F.1. First we note that the SDP 2 has $\ell$ smooth gradient for $\ell\leq poly(\rho,B)$ and has $\gamma$ lipschitz Hessian for $\gamma=poly(\rho,B)$ which we arrive at in Lemma E.2 and Lemma F.1. The proofs of Lemma E.2 and Lemma F.1 are technically involved and form the core hurdle in arriving at our proof.
Then by Theorem F.1 the iteration converges to an $(\epsilon^{\prime},\gamma^{2})$-SOSP Definition Definition in no more than $\tilde{O}(\frac{1}{\epsilon^{\prime 2}})$ iterations with probability $1-\delta$. Now that we’ve established that the basic gradient iteration converges to approximate SOSP, we need to then prove that the approximate SOSP are approximate global optimum. We achieve this by using the result of bhojanapalli2018smoothed lemma 3 which we restate in Lemma C.2 for convenience.
@@
Let the global optimum of SDP 2 be denoted $\tilde{\mathbb{E}}_{\tilde{\mu}}[\tilde{\mathbf{x}}]$ with a cholesky decomposition $\tilde{\mathbf{v}}$. Let $\hat{\mathbf{v}}$ be the set of vectors outputted by Algorithm 3 with associated pseudoexpectation $\tilde{\mathbb{E}}_{\hat{\mu}}[\hat{\mathbf{x}}]$. Then, we can bound
@@
Here the first equality is by definition, and the inequality is by the convexity of $\mathcal{H}_{\rho}$. Moving on, we use the fact that the min eigenvalue of the hessian of overparameterized loss $\nabla^{2}\mathcal{L}_{\rho}(\hat{\mathbf{v}})\succeq-\gamma\sqrt{\epsilon^{\prime}}$ implies the min eigenvalue of the gradient of the convex loss $\lambda_{\text{min}}(\nabla\mathcal{H}_{\rho}(\tilde{\mathbb{E}}[\hat{\mathbf{x}}]))\geq-\gamma\sqrt{\epsilon^{\prime}}$. This fact is invoked in bhojanapalli2018smoothed lemma 3 which we adapt to our setting in Lemma C.2. Subsequently, we adapt the lines of their argument in Lemma C.2 most relevant to our analysis which we detail here for the sake of completeness.
@@
Here the first inequality follows by a standard inequality of frobenius inner product, the second inequality follows by Cauchy-Schwarz, the third inequality follows by the $(\epsilon^{\prime},\gamma^{2})$-SOSP conditions on both the min eigenvalue of the hessian and the norm of the gradient, the final two inequalities follow from knowing the main diagonal of $\tilde{\mathbb{E}}[\hat{\mathbf{x}}]$ is the identity and that every vector in $\tilde{\mathbf{v}}$ is a unit vector up to inverse polynomial error $poly(\rho^{-1},2^{k})$. For this last point see the proof in Lemma E.1. Therefore if we set $\epsilon^{\prime}=poly(\epsilon,2^{-k})$ we arrive at any $\epsilon$ error. Therefore we have established our estimate $\hat{v}$ is approximates the global optimum of the quadratically penalized objective i.e $\mathcal{H}_{\rho}(\tilde{\mathbb{E}}[\hat{\mathbf{x}}])-\mathcal{H}_{\rho}(\tilde{\mathbb{E}}[\tilde{\mathbf{x}}])\leq\epsilon$. To finish our proof, we have to bound the distance between the global optimum of the quadratically penalized objective $\mathcal{H}_{\rho}(\tilde{\mathbb{E}}[\tilde{\mathbf{x}}])$ and SDP$(\Lambda)$ the optimum of SDP 2. This is established for $\rho$ a sufficiently large $poly(\epsilon^{-1},2^{k})$ in Lemma E.1. This concludes our proof that the iterates of Algorithm 3 converge to the solution of SDP 2. ∎
In our proof above, the bound on approximate SOSP being approximate global optimum is built on the result of bhojanapalli2018smoothed which we rephrase for our setting below.
@@
[bhojanapalli2018smoothed lemma 3 rephrased] Let $\mathcal{L}_{\rho}(\cdot)$ be defined as in equation 63 and let $\mathcal{H}_{\rho}(\cdot)$ be defined as in equation 69. Let $U\in\mathbb{R}^{\Phi\times\Phi}$ be an $(\epsilon,\gamma)$-SOSP of $\mathcal{L}_{\rho}(\cdot)$, then
@@
The following Lemma E.1 establishes that for a sufficiently large penalty parameter $\rho=poly(\epsilon^{-1},2^{k})$ the optimum of the penalized problem and the exact solution to SDP 2 are close.
@@
Let $\Lambda$ be a Max-k-CSP instance, and let SDP$(\Lambda)$ be the optimum of SDP 2. Let $\mathcal{L}_{\rho}(\mathbf{v})$ be the quadratically penalized objective
@@
For an objective matrix $C$ and constraint matrices $\{A_{i}\}_{i\in\mathcal{F}}$ in some constraint set $\mathcal{F}$. We will invoke specific properties of SDP 2 to enable our analysis. First we define the penalized objective in this generic form
@@
We upper bound the RHS using the robustness theorem of RSanycsp09 restated in the appendix Theorem F.2 which states that an SDP solution that violates the constraints by a small perturbation changes the objective by a small amount. Thus we obtain,
@@
SDP 3 SDP Vector Formulation for Max-k-CSP General Alphabet (Equivalent to UGC optimal)
@@
C.1 Neural Certification Scheme